PIP-II report for May 11, 2018

 

Highlights:

 

Project Management (Marc Kaducak):

  • Held our first weekly meeting with the FNAL procurement department and the Technical Division materials group to improve communications between all parties. Provided a list of procurements >$100k to the DOE IPT and shared with the FNAL procurement department.
  • Finalized the assumptions for in-kind SRF and RF Power deliverables to be used in developing the resource loaded schedule (RLS).
  • Worked with OPSS on the structure and required content of our Basis of Estimate (BoE) documents.
  • Kicked off RLS development efforts with LLRF and Accelerator Complex Upgrades.

 

Technical Integration (Arkadiy Klebaner):

  • Completed Design Readiness Plan and Design Deliverables Document for CD-2.
  • Requirements Verification work is proceeding; FRS completion is at 40%. Completed generic safety requirements statement for inclusion into all FRS’s.
  • Interface development is advancing – 45% completion.
  • Started work on the consolidated Preliminary Design Review plan.

 

International Coordination (Steve Holmes & Lia Merminga):

  • 650 MHz/40 kW RF Amplifier TRS has been approved by DAE.
  • Updated list of Fermilab SPMs and DAE SPCs finalized.
  • Five DAE employees have been approved to visit FNAL for 18 months.

 

SRF & Cryo Systems (Allan Rowe & Genfa Wu):

HWR

  • HWR CM Production
    • ANL finished the repair of the upper lid. Preparation for the Argonne internal safety review of the HWR CM assembly started.
    • The review of HWR Interface Specification Document (ICD) is ongoing.
    • Modified HWR CM, Cavity and Power Coupler FRSs in order to comply with the latest PIP-II CDR revision
  • HWR Integrated Cold Test with Power Coupler
    • 7 kW TOMCO solid-state amplifier is now connected to the building electrical network.
    • The Cold Test of the HWR #5 with power coupler is scheduled for next week.
    • The procedure of the HWR Power Coupler RF test and acceptance criteria have to be finalized.

SSR1

  • SSR1 Cavities, coupler and tuners
    • Setup for next cavities testing in the STC will include the installation of warm-end coupler and tuner active components (motor and piezo). Qualified components will be assigned to the specific cavity and stored along with kits for cryomodule assembly.
    • S111 failed qualification with HPC at STC due to strong FE.
    • S113 received 120C baking and is installed in the STC. Cooldown over the weekend. Testing next week.
    • S106 transported to ANL for light BCP and HPR, LPC preparation after flange maintenance.
    • Cavity S107 is prepared in the cleanroom at Lab2 to receive HPC next week.
    • Procedure for the string assembly will be ready within 2 weeks.
    • PO was issued to procure 6 vacuum-end units needed to complete the qualification of the cavities.
  • SSR1 Cryomodule
    • Current leads investigation is in progress. 2 out of 4 sets of current leads for the HWR were leak checked and they were found leaking at the interface between the polyolefin tube and copper wire. A meeting will be set next week to discuss possible modification for the HWR leads and to finalize the design of the SSR1 ones.
    • Requisition for thermal shields, flexible hoses and cooldown/warm are approved.
    • Design of lifting tooling: completed.

HB650

  • B9AS-RRCAT-301 was tested at FNAL, at 2 K, 1.8 K, and 1.4 K. Cavity displayed field emission beginning at ~12 MV/m, and quenched at 15 MV/m.
    • This cavity will be re-washed (I.e. HPR) and retested.
  • Collaboration meeting held.
  • Couplers received and storage tooling being designed.
  • RF Coupler development meeting held. Acceptance and verification criteria discussed.

Cryogenics Plant

  • Finished reviewing Cryoplant Building TRS and will meet with CF Manager in the week of May 14 to discuss details.
  • Trip/visit to ESS booked for Week of May 21 to investigate ESS Cryoplant and integration.
  • Preparing for trip to ESS on the week of May 21st. Plan to have discussions regarding the ESS CDS which is very similar to the PIP-II CDS.
  • Continued work on ICD development.

 

Accelerator Systems (Paul Derwent):

High Power RF

  • 650 MHz 40kW RF Power Amplifier TRS has been approved by Indian collaborators.  It will be circulated for FNAL approval.
  • MDB Coupler Test – working on 3-D model for the circulator stand.  Detailed 2-D drawing and the list of transmission line parts we need to order exist.

Low Level RF

  • Purchase requisitions have been put through for LLRF hardware needed for PIP2IT.
  • PIP2IT beam studies with RFQ system setup were performed.

Instrumentation

  • The prototype wire scanner was put under vacuum and aligned early in the week. On Friday, first beam profile measurements at PIP2IT were made. The system performed quite well, a few things are required to make it operational. The L3 manager (Vic) would like to express thanks to everyone who has worked on this system.

Prototype Wire Scanner installed in the beamline at PIP2IT.

Vacuum

  • On 5/4, the modified right side plate on RFQ Module3 pumping port was re-installed and RFQ vacuum was restored. Helium was injected around the seal on 5/4 and 5/8. RGA data was analyzed along with ACNET data, no bursts were found from this pumping port after modification.
  • Draft PDR Vacuum system outline was submitted to the editors.

 

Linac Installation & Commissioning (Fernanda G. Garcia):

PIP2IT:

  • During last week shutdown, the main work was related to understand the RFQ vacuum spikes.
    • One of the blank-off was removed and machined to flattened out the vacuum side, especially around a weld region.
    • Cover was reinstalled and preliminary vacuum and RGA tests showed no helium peaks concurrent with large vacuum bursts.

Test Infrastructure:

  • PIP2IT Cryo transfer line support columns were installed, and first sections of external piping are welded and awaiting installation by rigging crew (see pictures below).

 

Installation:

  • Continue making progress on PIP-II 3D CAD model and agreement was reached on the preliminary format for PIP-II lattice data output. Optics team will be updating scripts to translate PIP-II information to the desired format, followed by a test run of the import into CAD program.
  • Began preparation of the “blocks” that will be populated into the 3D CAD model by the lattice import script.  Focusing on cryomodules and magnets to start with.

 

Conventional Facilities (Steve Dixon):

  • GPP Coordination Meeting – Met with FESS to understand the impact and timing of the candidate GPP projects in order to coordinate with the ongoing design of the Site Prep package.